OpenStudy (anonymous):

If you were to use the substitution method to solve the following system, choose the new equation after the expression equivalent to x from the first equation is substituted into the second equation. -x + 5y = 1 2x + 4y = -4 Choose one answer. a. 2(5y - 1) + 4y = -4 b. 2(-5y + 1) + 4y = -4 c. 2x + 4(-5y + 1) = -4 d. 2x + 4(5y - 1) = -4

OpenStudy (ja1):

XD i actually just learned this lol

OpenStudy (anonymous):

so you will be putting the first equation into the second. in this one you are solving for x for the first equation, do you know how to do this

OpenStudy (anonymous):

can you just show your work, for the whole thing

OpenStudy (anonymous):

so the first equation is -x +5y = 1 You want to solve for x, what that means is getting x alone on one side of the equation. you can do this many ways, i will add x to each side to get -x + x + 5y = 1 +x the x's on the left side cancel to zero 5y = 1 + x subtract 1 from each side 5y - 1 = 1 - 1 + x 5y - 1 = x :since 1 - 1 = 0 now that you have x = 5y - 1 you can put it into the second equation 2x + 4y = -4 :just replace x with 5y -1 since you know they are equal they are the same thing 2( 5y-1) + 4y = -4

OpenStudy (anonymous):

can you help me with another

OpenStudy (anonymous):

sure thing

OpenStudy (anonymous):

If you were to use the substitution method to solve the following system, choose the new system of equations that would result if z was isolated in the first equation. 2x - 5y + z = 15 4x - y + 3z = 15 3x - 2y + 2z = 13

OpenStudy (anonymous):

this time i want you to get z on its own, 2x - 5y + z = 15 you want z on one side and the rest on the other. Once you get that i will continue

OpenStudy (anonymous):

alright one sec

OpenStudy (anonymous):

is it z=15+5y

OpenStudy (anonymous):

z=15+5y-2x

OpenStudy (anonymous):

yes it is.

OpenStudy (anonymous):

Now that you have z, z=15+5y-2x, you just replace that z into both of the equations second equation: 4x - y + 3z = 15 so putting z in: 4x - y + 3(15+5y-2x) = 15 third equation: 3x - 2y + 2z = 13 with z substituted: 3x - 2y + 2(15+5y-2x) = 13

OpenStudy (anonymous):

Now that more difficult part is to simplify both of these all down to just one x term, one y term, and one constant each. But once you do it would be just like the first question with 2 variables

OpenStudy (anonymous):

now the* more difficult part

OpenStudy (anonymous):

do you need to see how the simplification would go?

OpenStudy (anonymous):

no it's ok

OpenStudy (anonymous):

i follow you